OpenStudy (anonymous):

Given f(x) = x^6 + 1 (a) Find all six complex roots f(x) = 0 (b) Find the angle between successive roots It's fine if you can't get (b), but (a) is what I really need now.

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

The roots will be all the sixth roots of -1. You have \[\begin{align*} x^6+1&=0\\ x^6&=-1\\ x&=(-1)^{1/6}\\ x&=\large\left(e^{i\pi}\right)^{1/6}\\ x&=\Large e^{i\left(\frac{\pi}{6}+\frac{2\pi k}{6}\right)} \end{align*}\] where \(k=0,1,...,5\).

OpenStudy (anonymous):

we never learned e ^i pi what is that?

OpenStudy (anonymous):

It's a part of Euler's formula, \[\large e^{i\theta}=\cos\theta+i\sin\theta\]

OpenStudy (anonymous):

ohh in our class we use z

OpenStudy (anonymous):

where did u get 30 degrees from

OpenStudy (anonymous):

pi/6

OpenStudy (anonymous):

wait oh sorry "z" in our class is =r(cos theta + isin theta)

OpenStudy (anonymous):

The polar form of \(-1\) is \(e^{i\pi}\), or \(\cos\pi+i\sin\pi=-1+0i=-1\). Raising this to the 1/6 power, DeMoivre's theorem says that \[\large\left(re^{i\theta}\right)^n=r^ne^{in\theta}\] or \[\large\left(r(\cos\theta+i\sin\theta)\right)^n=r^n(\cos n\theta+i\sin n\theta)\] What this means is that \[\large \left(e^{i\pi}\right)^{1/6}=e^{i\pi/6}\] where the angle is in radians, or 30 degrees (relative to the positive real axis).

OpenStudy (anonymous):

can you use arctan(b/a) to find theta? Because that is the only way I learned how to do it

OpenStudy (anonymous):

Have you not learned DeMoivre's theorem?

OpenStudy (anonymous):

\[z ^{n}=[r(\cos \theta + isin \theta)]^{n}=r^n(\cos n \theta + i \sin n \theta)\]

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

thats how we learned it

OpenStudy (anonymous):

Okay, well \(\theta\) is the argument/angle of the original number (-1 in this case with angle \(\pi\), or 180 deg). When you raise the polar form to 1/6, you divide whatever this angle is by 6 (so that you have \(\dfrac{\pi}{6}\), or 30 deg). You can use the inverse tangent approach to find the angle of -1, but it won't give you the angle of the principal sixth root of -1 unless you divide by 6.

OpenStudy (anonymous):

so the n is 1/6 and when you multiply that by theta in the formula you get 30?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

then how would you find all 6 roots? +2kpi?

OpenStudy (anonymous):

\(+\dfrac{2k\pi}{n}\), yes

OpenStudy (anonymous):

This should give you an idea of how to do part (b). Each angle between successive roots is the same.

OpenStudy (anonymous):

so when k=1, is the theta 5 degrees?

OpenStudy (anonymous):

No, for \(k=0\) you get your principal root, which has the angle \[\frac{\pi}{6}+\frac{2\times0 \pi}{6}\] For \(k=1\), you get \[\frac{\pi}{6}+\frac{2\times1 \pi}{6}=\frac{3\pi}{6}=\frac{\pi}{2}\] and so on.

OpenStudy (anonymous):

do you know (b)? because if you don't that is completely fine

OpenStudy (anonymous):

any ideas?

OpenStudy (anonymous):

What's the angle between the first and second roots?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

There you go.

OpenStudy (anonymous):

thats it?

OpenStudy (anonymous):

Yes, the agnle between successive roots is the same for all successive roots. Consider the Argand diagram: